OpenStudy (anonymous):

The equation below represents Function A and the graph represents Function B: Function A f(x) = -2x + 1 Function B graph of line going through ordered pairs negative 1, negative 5 and 2, 1 and 3, 3 Which equation best compares the slopes of the two functions? Slope of Function B = 2 × Slope of Function A. Slope of Function A = Slope of Function B Slope of Function A = 2 × Slope of Function B Slope of Function B = - Slope of Function A

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

The first five terms of a sequence are shown below: 4, 7, 10, 13, 16 If the nth term of this sequence is represented by f(n), which of the following functions best represents this sequence? f(n) = 3n - 1; n ≥ 1 f(n) = 3n + 1; n ≥ 1 f(n) = 4n - 1; n ≥ 1 f(n) = 4n + 1; n ≥ 1

OpenStudy (anonymous):

The description below represents Function A and the table represents Function B: Function A The function is 8 more than 3 times x. Function B x y -1 2 0 5 1 8 Which statement is correct about the slope and intercept of the two functions? Their slopes are equal, but y-intercepts are not equal. Their slopes are not equal, but y-intercepts are equal. Both slopes and y-intercepts are equal. Neither slopes nor y-intercepts are equal.

OpenStudy (anonymous):

That was all @ganeshie8

ganeshie8 (ganeshie8):

@jadee27p

ganeshie8 (ganeshie8):

For the first question, can you find the `slope` and `y intercept` of both the given functions ?

ganeshie8 (ganeshie8):

``` Function A f(x) = -2x + 1 ``` slope = ? y intercept = ?